1. Write a balanced chemical equation, including states of matter, for the combustion of gaseous benzene, C6H6.
2.Write a balanced chemical equation, including states of matter, for the combustion of ethylene gas, C2H4.
3. How many grams of CaF2 would be needed to produce 1.01 moles of F2?
CaF₂ (s) + H₂SO₄ (l) → 2HF(l) + CaSO₄ (s)
2HF (l) → F₂ (g)+ H₂ (g)
1
Expert's answer
2019-11-12T12:02:17-0500
"2C_6H_6(g)+ 15O_2{(g)} = 12CO_2(g) + 6H_2O(g)"
  1. At high temperatures characteristic of the conditions of combustion reactions, the water formed is most likely gaseous.
"C_2H_4(g) + 3O_2(g) = 2CO_2(g) + 2H_2O(g)"
  1. According to the reaction scheme:
"n(CaF_2) = n(F_2) = 1.01 mol"

"m(CaF_2) = n*M = 1.01mol*78.07g\/mol = 78.85g"